# Understanding Before Recommendation: Semantic Aspect-Aware Review Exploitation via Large Language Models

### v1

@@ 5. Conclusion
Our work proposes a novel approach that extracts semantic aspects from user reviews and utilizes them to enhance recommendation performance. Specifically, we devise a chain-based prompting approach with LLMs to extract the semantic aspect-aware reviews, which are then utilized to identify semantic aspect-aware user-item interactions. To improve the recommendation performance and interpretability, we propose a semantic aspect-based graph convolution network (SAGCN) that performs graph convolution operations on multiple semantic aspect-based graphs. We conducted extensive experiments on three benchmark datasets to evaluate the performance of our SAGCN model. Our experimental results demonstrate that SAGCN outperforms state-of-the-art GCN-based recommendation models. The experiment analysis exhibits the interpretability of our approach.
In summary, our work emphasizes the importance of analyzing user reviews by creating structured reviews to understand users before recommendations. On the other hand, we investigate a new approach to improving recommendation performance by utilizing LLMs.

### v2

@@ 5. Conclusion and Future Work
Our paper proposes a novel approach that extracts semantic aspects from user reviews and utilizes them to enhance recommendation accuracy and interpretability. Specifically, we introduce a chain-based prompting strategy with LLMs to extract the semantic aspect-aware reviews, which are then utilized to identify semantic aspect-aware user-item interactions. To exploit these interactions for representation learning of users and items, we propose a semantic aspect-based graph convolution network (SAGCN) that performs graph convolution operations on multiple semantic aspect-based graphs. We conducted extensive experiments on four benchmark datasets to evaluate the performance of our SAGCN model. Our experimental results demonstrate that SAGCN outperforms the pure LLM-based method and the state-of-the-art GCN-based recommendation models. The experiment analysis exhibits the interpretability of our approach. In summary, our work emphasizes the importance of analyzing user reviews by creating structured reviews to understand users before recommendations. On the other hand, we investigate a new approach to improving recommendation performance by integrating the capability of LLMs with conventional recommendation technique.
This work is an initial attempt to combine the sentiment analysis capabilities of LLMs with conventional recommendation techniques. Our approach involves using a frozen LLM to identify aspect-aware reviews. However, accurately extracting semantic aspects from user reviews remains a challenging task. In addition, apart from the reviews provided by users, the interactions between users and items also include semantic aspects. We expect the potential of sentiment analysis to be further explored for a better understanding of user intent. Furthermore, it would be interesting to explore how to take advantage of these aspect-aware user-item interactions for representation learning.